# TeSent: A Benchmark Dataset for Fairness-aware Explainable Sentiment Classification in Telugu

### v1

@@ B.1. Gender Bias
To assess gender bias, 1320 counterfactual pairs are generated. This includes i) 440 neutral pairs created using the four templates from Type(b) with 100 random male-female named pairs stratified across Hindu, Muslim, and Christian names, plus 10 noun-phrase pairs, per template and ii) 440 positive and negative pairs each, stratified across the three religions, using the seven templates from Type(a): for each of 10 randomly chosen ¡emotion¿ word of a polarity, generate 44 pairs by randomly selecting a template and a random named pair or noun phrase pair.
B.2. Religion Bias
@@
To generate Religion-based Bias Dataset, we adapt a similar sampling procedure as discussed for Gender-based Bias Dataset, but discarding the noun phrases, as they do not contribute to any religion. We generate 1200 counterfactual pairs for each combination of the three religions (Hindu, Muslim, and Christian). This includes, i) 400 neutral pairs created using Type(b) templates with 100 random named pairs of a chosen religion combination stratified across female and male groups, per template and ii) 400 positive and negative pairs each, stratified across the two gender groups, using the seven templates from Type(a): for each 10 randomly chosen ¡emotion¿ word of a polarity, generate 40 pairs by randomly selecting a template and a random named pair.
Appendix C Pre-trained Models
C.1. mBERT
@@ Appendix C Pre-trained Models
mBERT (BERT-base-multilingual-cased) is based on Google’s BERT-base transformer model (with 12 Layers and 100$\sim$ 100∼ 100 million parameters) trained on Wikipedia texts in 104 languages, including Telugu for the Masked Language Modeling (MLM) and Next Sentence Prediction tasks [Devlin et al., 2019]. Although not specifically trained on Telugu alone, mBERT has demonstrated strong performance in sentiment classification tasks due to its shared multilingual representation [Hedderich et al., 2021]. For Telugu sentiment classification, mBERT supports cross-lingual transfer capabilities with acceptable performance even with limited data on Telugu. Its ability to generalize across languages makes it effective for multilingual applications[Kalyan et al., 2021], although it may not capture fine-grained Telugu-specific nuances as well as regionally tuned models. However, since it is not optimized for Telugu morphology or syntax, its performance may lag [Wu and Dredze, 2020] behind Telugu-specialized models such as IndicBERT, L3Cube-Telugu-BERT, in capturing language-specific nuances. Nonetheless, mBERT remains a powerful yet reliable baseline that is widely used in academic research, particularly in low-resource settings [Marreddy et al., 2022, Rajalakshmi et al., 2023, Park et al., , Marreddy et al., 2021, Duggenpudi et al., 2022].
C.2. XLM-R
@@
XLM-RoBERTa (XLM-R) is a general-purpose multilingual transformer model developed by Facebook AI, specifically designed to improve cross-lingual understanding by training on a massive multilingual corpus (2.5TB of filtered “Common Crawl” data across 100+ languages, including Telugu) for MLM task (with no NSP as used in mBERT) [Conneau et al., 2019]. As it purely focuses on MLM tasks, it provides better contextual modeling and transfer learning than mBERT, having improved downstream performance [Hedderich et al., 2021]. Therefore, XLM-R can lend good performance for Telugu sentiment analysis, particularly when fine-tuned with local data for benchmarking[Kulkarni et al., 2021, Joshi, 2022]. However, Telugu-specific models like MuRIL or L3Cube-Telugu-BERT may offer better cultural and linguistic alignment [Das et al., 2022, Rajalakshmi et al., 2023].
C.3. MuRIL
@@
MuRIL (Multilingual Representations for Indian Languages) is also a transformer-based BERT model, specifically designed to support 17+ Indian languages, including Telugu and English [Khanuja et al., 2021]. Unlike mBERT, MuRIL is pre-trained on a large corpus of Telugu sentences from web, religious script, news data, etc., for the Masked Language Modeling (MLM) and Translation Language Modeling (TLM) tasks, offering better understanding of Telugu morphology and syntax [Joshi, 2022]. As the pre-training data favors informal texts from the web, MuRIL can be less effective for formal and classical Telugu NLP tasks. However, MuRIL is a strong candidate for Telugu sentiment classification, especially when analyzing informal, social media, or conversational data, making it superior to general multilingual models like mBERT and XLM-R [Das et al., 2022, Rajalakshmi et al., 2023].
C.4. IndicBERT
@@
IndicBERT (ai4bharat/indicBERTv2-MLM-only) is a multilingual BERT-like model developed by AI4Bharat, trained on OSCAR and AI4Bharat curated corpora of 12 Indian languages (including Telugu and English) for MLM task [Joshi, 2022]. It is well-suited for monolingual Telugu NLP tasks rather than cross-lingual ones and do not support code-mixed data. For Telugu sentiment classification, it provides language-aware tokenization, clean embeddings, and faster training [Marreddy et al., 2022, Rajalakshmi et al., 2023, Duggenpudi et al., 2022].
C.5. Te-BERT
@@
Telugu-BERT (L3Cube-Telugu-BERT) is a transformer-based BERT model pre-trained specifically on Telugu text (Telugu OSCAR, Wikipedia, news) by the L3Cube Pune research group for MLM task [Joshi, 2022]. Since Telugu-BERT is tailored for Telugu, it excels in capturing the vocabulary, syntax, and semantics of the language, thus able to recognize nuanced expressions, idioms, and sentiments that are often poorly represented in multilingual models like mBERT and XLM-R. Therefore, it is ideal for researchers working on pure Telugu text analysis with sufficient labeled data for fine-tuning.
Appendix D Hyper-parameter Tuning
@@
We split the dataset into 80% training, 10% validation, and 10% test sets using stratified sampling. On the validation set, we applied grid search hyperparameter tuning across all selected models. For each model, we explored a range of batch sizes (16, 32, 64) and learning rates (LR) from 1e-5 to 5e-5. For training with rationale, we varied the regularization parameter $\lambda$ from 0.1 to 0.7. After extensive tuning, the best performance was observed with $\lambda=0.6$, which consistently delivered optimal results across all models. Finally, for both types of training, we used batch size as 64, 2e-5 as LR, and the Adam optimizer, and obtained consistently strong results with 4 epochs.
Appendix E Annotation Interface
@@
Figure 6. Annotation Interface
Appendix F Telugu Topics Overview
[…]
@@ Appendix F Telugu Topics Overview
Religion: ”Role of religion in Telugu society”, ”Telugu religious beliefs and practices”, ”Influence of Telugu faith in modern politics”
@@
Education: ”Telugu education system reform”, ”Online Telugu learning platforms”, ”Future of education in Telugu medium schools”
@@
Personality: ”Telugu personality development programs”, ”Analysis of Telugu personality traits”, ”Psychology of Telugu personalities”
@@
Food & Cooking: ”Popular Telugu recipes and dishes”, ”Modern trends in Telugu cooking”, ”Healthy eating in Telugu cuisine”
@@
Shopping: ”Telugu online shopping behaviour”, ”Telugu consumer preferences and habits”, ”Black Friday deals in Telugu markets”
@@
Crime: ”Telugu crime rate analysis”, ”Notable Telugu true crime cases”, ”Telugu criminal justice system overview”
@@
Law & Justice: ”Telugu legal system reform efforts”, ”Access to justice in Telugu communities”, ”High-profile Telugu court case summaries”
@@
Trade: ”Telugu international trade discussions”, ”Impact of trade wars on Telugu businesses”, ”Telugu trade policy analysis”
@@
Entertainment: ”Latest Telugu entertainment updates”, ”Trends in Telugu Tollywood industry”, ”Telugu movie reviews and ratings”
@@
Sports: ”Highlights of Telugu sports achievements”, ”Telugu athletes in Olympics history”, ”Debates in Telugu sports media”
@@
Technology: ”Innovations in Telugu tech startups”, ”Role of AI in Telugu applications”, ”Future of Telugu digital technologies”
@@
People: ”Inspiring Telugu personal stories”, ”Influential Telugu people in history”, ”Interviews with Telugu celebrities”
@@
Science: ”Scientific breakthroughs by Telugu researchers”, ”Latest Telugu science fair projects”, ”Space exploration in Telugu media”
@@
Government: ”Telugu government policy updates”, ”Public administration in Telugu regions”, ”Governance challenges in Telugu states”
@@
Health: ”Telugu health awareness campaigns”, ”Mental health in Telugu society”, ”Overview of Telugu healthcare systems”
@@
Product: ”Telugu tech product reviews”, ”Best Telugu consumer electronics”, ”Top Telugu products of 2024”
@@
International: ”International Telugu diaspora news”, ”Global events affecting Telugu communities”, ”Telugu perspectives on international relations”
@@
Election: ”Telugu election coverage highlights”, ”Voter turnout trends in Telugu states”, ”Analysis of Telugu election results”
@@
Service: ”Customer service trends in Telugu markets”, ”Best service industries for Telugu users”, ”Innovations in Telugu public service delivery”
@@
Politics: ”Telugu political debates and opinions”, ”Current affairs in Telugu politics”, ”Telugu party manifestos and analysis”
@@
Job: ”Telugu job market opportunities”, ”Career guidance for Telugu graduates”, ”Getting hired in Telugu job portals”
@@
Social Media: ”Telugu social media influencers”, ”Trends on Telugu social platforms”, ”Impact of social media on Telugu youth”
@@
News: ”Breaking Telugu news stories”, ”World news from Telugu perspective”, ”Top Telugu news highlights”
@@
Environment: ”Telugu discussions on climate change”, ”Sustainability efforts in Telugu regions”, ”Environmental issues in Telugu states”
@@
Accident: ”Telugu accident news reports”, ”Road safety tips for Telugu drivers”, ”Prevention of accidents in Telugu regions”
@@
Transportation: ”Public transport in Telugu cities”, ”Rise of electric vehicles in Telugu states”, ”Future of Telugu transportation systems”
@@
Economics: ”Telugu economic development trends”, ”State of Telugu regional economies”, ”Economic theory in Telugu discourse”
@@
Stock Market: ”Telugu stock market tips and strategies”, ”Investing trends among Telugu traders”, ”Telugu market news and updates”
@@
Others: ”Trending topics in Telugu discussions”, ”Popular Telugu YouTube content”, ”Miscellaneous Telugu news highlights”

### v2

@@ I.1 Gender Bias
To assess gender bias, 1320 counterfactual pairs are generated. This includes i) 440 neutral pairs created using the four templates from Type(b) with 100 random male-female named pairs stratified across Hindu, Muslim, and Christian names, plus 10 noun-phrase pairs, per template and ii) 440 positive and negative pairs each, stratified across the three religions, using the seven templates from Type(a): for each of 10 randomly chosen <emotion> word of a polarity, generate 44 pairs by randomly selecting a template and a random named pair or noun phrase pair.
I.2 Religion Bias
@@
To generate Religion-based Bias Dataset, we adapt a similar sampling procedure as discussed for Gender-based Bias Dataset, but discarding the noun phrases, as they do not contribute to any religion. We generate 1200 counterfactual pairs for each combination of the three religions (Hindu, Muslim, and Christian). This includes, i) 400 neutral pairs created using Type(b) templates with 100 random named pairs of a chosen religion combination stratified across female and male groups, per template and ii) 400 positive and negative pairs each, stratified across the two gender groups, using the seven templates from Type(a): for each 10 randomly chosen <emotion> word of a polarity, generate 40 pairs by randomly selecting a template and a random named pair.
Appendix J Pre-trained Models
J.1 mBERT
@@
mBERT (BERT-base-multilingual-cased) is based on Google’s BERT-base transformer model (with 12 Layers and $\sim 100$ million parameters) trained on Wikipedia texts in 104 languages, including Telugu for the Masked Language Modeling (MLM) and Next Sentence Prediction tasks Devlin et al. (2019). Although not specifically trained on Telugu alone, mBERT has demonstrated strong performance in sentiment classification tasks due to its shared multilingual representation Hedderich et al. (2021). For Telugu sentiment classification, mBERT supports cross-lingual transfer capabilities with acceptable performance even with limited data on Telugu. Its ability to generalize across languages makes it effective for multilingual applicationsKalyan et al. (2021), although it may not capture fine-grained Telugu-specific nuances as well as regionally tuned models. However, since it is not optimized for Telugu morphology or syntax, its performance may lag Wu and Dredze (2020) behind Telugu-specialized models such as IndicBERT, L3Cube-Telugu-BERT, in capturing language-specific nuances. Nonetheless, mBERT remains a powerful yet reliable baseline that is widely used in academic research, particularly in low-resource settings Marreddy et al. (2022); Rajalakshmi et al. (2023); Park et al. ; Marreddy et al. (2021); Duggenpudi et al. (2022).
J.2 XLM-R
@@
XLM-RoBERTa (XLM-R) is a general-purpose multilingual transformer model developed by Facebook AI, specifically designed to improve cross-lingual understanding by training on a massive multilingual corpus (2.5TB of filtered “Common Crawl” data across 100+ languages, including Telugu) for MLM task (with no NSP as used in mBERT) Conneau et al. (2019). As it purely focuses on MLM tasks, it provides better contextual modeling and transfer learning than mBERT, having improved downstream performance Hedderich et al. (2021). Therefore, XLM-R can lend good performance for Telugu sentiment analysis, particularly when fine-tuned with local data for benchmarkingKulkarni et al. (2021); Joshi (2022). However, Telugu-specific models like MuRIL or L3Cube-Telugu-BERT may offer better cultural and linguistic alignment Das et al. (2022); Rajalakshmi et al. (2023).
J.3 MuRIL
@@
MuRIL (Multilingual Representations for Indian Languages) is also a transformer-based BERT model, specifically designed to support 17+ Indian languages, including Telugu and English Khanuja et al. (2021). Unlike mBERT, MuRIL is pre-trained on a large corpus of Telugu sentences from web, religious script, news data, etc., for the Masked Language Modeling (MLM) and Translation Language Modeling (TLM) tasks, offering better understanding of Telugu morphology and syntax Joshi (2022). As the pre-training data favors informal texts from the web, MuRIL can be less effective for formal and classical Telugu NLP tasks. However, MuRIL is a strong candidate for Telugu sentiment classification, especially when analyzing informal, social media, or conversational data, making it superior to general multilingual models like mBERT and XLM-R Das et al. (2022); Rajalakshmi et al. (2023).
J.4 IndicBERT
@@
IndicBERT (ai4bharat/indicBERTv2-MLM-only) is a multilingual BERT-like model developed by AI4Bharat, trained on OSCAR and AI4Bharat curated corpora of 12 Indian languages (including Telugu and English) for MLM task Joshi (2022). It is well-suited for monolingual Telugu NLP tasks rather than cross-lingual ones and do not support code-mixed data. For Telugu sentiment classification, it provides language-aware tokenization, clean embeddings, and faster training Marreddy et al. (2022); Rajalakshmi et al. (2023); Duggenpudi et al. (2022).
J.5 Te-BERT
@@
Telugu-BERT (L3Cube-Telugu-BERT) is a transformer-based BERT model pre-trained specifically on Telugu text (Telugu OSCAR, Wikipedia, news) by the L3Cube Pune research group for MLM task Joshi (2022). Since Telugu-BERT is tailored for Telugu, it excels in capturing the vocabulary, syntax, and semantics of the language, thus able to recognize nuanced expressions, idioms, and sentiments that are often poorly represented in multilingual models like mBERT and XLM-R. Therefore, it is ideal for researchers working on pure Telugu text analysis with sufficient labeled data for fine-tuning.
Appendix K Hyper-parameter Tuning
@@
We split the dataset into 80% training, 10% validation, and 10% test sets using stratified sampling. On the validation set, we applied grid search hyperparameter tuning across all selected models. For each model, we explored a range of batch sizes (16, 32, 64) and learning rates (LR) from 1e-5 to 5e-5. For training with rationale, we varied the regularization parameter $\lambda$ from 0.1 to 0.7. After extensive tuning, the best performance was observed with $\lambda=0.7$, which consistently delivered optimal results across all models. Finally, for both types of training, we used batch size as 64, 2e-5 as LR, and the Adam optimizer, and obtained consistently strong results with 4 epochs. We used a standard Python 3 Google Colab (Pro) runtime with one L4 GPU for all experiments, including model training and explanation generation, over a total duration of approximately 14 hours.
Appendix L Annotation Interface
@@
Figure 6: Annotation Interface
Appendix M Telugu Topics Overview
For brevity, we are listing the top 3 search phrases for each category, using which we collected $90\%$ of the comments.
@@
Religion: "Role of religion in Telugu society", "Telugu religious beliefs and practices", "Influence of Telugu faith in modern politics"
@@
Education: "Telugu education system reform", "Online Telugu learning platforms", "Future of education in Telugu medium schools"
@@
Personality: "Telugu personality development programs", "Analysis of Telugu personality traits", "Psychology of Telugu personalities"
@@
Food & Cooking: "Popular Telugu recipes and dishes", "Modern trends in Telugu cooking", "Healthy eating in Telugu cuisine"
@@
Shopping: "Telugu online shopping behaviour", "Telugu consumer preferences and habits", "Black Friday deals in Telugu markets"
@@
Crime: "Telugu crime rate analysis", "Notable Telugu true crime cases", "Telugu criminal justice system overview"
@@
Law & Justice: "Telugu legal system reform efforts", "Access to justice in Telugu communities", "High-profile Telugu court case summaries"
@@
Trade: "Telugu international trade discussions", "Impact of trade wars on Telugu businesses", "Telugu trade policy analysis"
@@
Entertainment: "Latest Telugu entertainment updates", "Trends in Telugu Tollywood industry", "Telugu movie reviews and ratings"
@@
Sports: "Highlights of Telugu sports achievements", "Telugu athletes in Olympics history", "Debates in Telugu sports media"
@@
Technology: "Innovations in Telugu tech startups", "Role of AI in Telugu applications", "Future of Telugu digital technologies"
@@
People: "Inspiring Telugu personal stories", "Influential Telugu people in history", "Interviews with Telugu celebrities"
@@
Science: "Scientific breakthroughs by Telugu researchers", "Latest Telugu science fair projects", "Space exploration in Telugu media"
@@
Government: "Telugu government policy updates", "Public administration in Telugu regions", "Governance challenges in Telugu states"
@@
Health: "Telugu health awareness campaigns", "Mental health in Telugu society", "Overview of Telugu healthcare systems"
@@
Product: "Telugu tech product reviews", "Best Telugu consumer electronics", "Top Telugu products of 2024"
@@
International: "International Telugu diaspora news", "Global events affecting Telugu communities", "Telugu perspectives on international relations"
@@
Election: "Telugu election coverage highlights", "Voter turnout trends in Telugu states", "Analysis of Telugu election results"
@@
Service: "Customer service trends in Telugu markets", "Best service industries for Telugu users", "Innovations in Telugu public service delivery"
@@
Politics: "Telugu political debates and opinions", "Current affairs in Telugu politics", "Telugu party manifestos and analysis"
@@
Job: "Telugu job market opportunities", "Career guidance for Telugu graduates", "Getting hired in Telugu job portals"
@@
Social Media: "Telugu social media influencers", "Trends on Telugu social platforms", "Impact of social media on Telugu youth"
@@
News: "Breaking Telugu news stories", "World news from Telugu perspective", "Top Telugu news highlights"
@@
Environment: "Telugu discussions on climate change", "Sustainability efforts in Telugu regions", "Environmental issues in Telugu states"
@@
Accident: "Telugu accident news reports", "Road safety tips for Telugu drivers", "Prevention of accidents in Telugu regions"
@@
Transportation: "Public transport in Telugu cities", "Rise of electric vehicles in Telugu states", "Future of Telugu transportation systems"
@@
Economics: "Telugu economic development trends", "State of Telugu regional economies", "Economic theory in Telugu discourse"
@@
Stock Market: "Telugu stock market tips and strategies", "Investing trends among Telugu traders", "Telugu market news and updates"
@@
Others: "Trending topics in Telugu discussions", "Popular Telugu YouTube content", "Miscellaneous Telugu news highlights"